OpenStudy (anonymous):

write the first 6 terms of the arithmetic sequences with the givin term a(1) and common differences d a1=200, d=10 (the one is a small lower on the bottom right cornor like a exponent)

OpenStudy (anonymous):

start with: a1 = 200; a2 = a1 + 10; a3 = a2 + 10; ... just keep adding 10 to the previous term to get the next term.